OpenStudy (anonymous):

A bee with velocity vector r'(t) starts out at the origin at t=0 and flies around for T seconds. Where is the bee located at time T if ∫(0-T) r'(u)du = 0? Also, what does the quantity ∫(0-T) ||r'(u)||du represent?

OpenStudy (anonymous):

Please help

OpenStudy (amistre64):

what does (0-T) represent?

OpenStudy (anonymous):

integral from 0 to T

OpenStudy (anonymous):

sorry it's the only way I know how to type it

OpenStudy (amistre64):

so those are the limits, ok

OpenStudy (amistre64):

do you recall the definition of: \[\int_{a}^{b}f(x)~dx=??\]

OpenStudy (amistre64):

the fundamental thrm of calculus is also what it is called

OpenStudy (anonymous):

It's F(b) - F(a)

OpenStudy (amistre64):

correct, so if F(b)-F(a) = 0, then either a=b, or F has the same value at multiple input levels similar to a periodic function.

OpenStudy (amistre64):

for example, cos(pi) - cos(0) = 0 since cos(pi)=cos(0)

OpenStudy (amistre64):

but the question is, what does taking the integral of a function tell us?

OpenStudy (anonymous):

it can tell you what the area under a curve is

OpenStudy (amistre64):

and my example is spose to read cos(2pi) = cos(0), but the site is lagging for me

OpenStudy (anonymous):

it's ok

OpenStudy (amistre64):

the area under a curve, yes ... but that reflects displacement

OpenStudy (amistre64):

ever wonder why in some cases we get a negative area? and in other cases we dont?

OpenStudy (anonymous):

Yes because the integral of velocity is displacement

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

negative displacement ?

OpenStudy (amistre64):

when we want to find the area between 2 curves, area is a positive value always ... its an absolute value. And in those cases we have to concern ourselves with the places that one curve crosses over the other.

OpenStudy (amistre64):

displacement on the other hand reflects is a vector, it has magnitude and direction.

OpenStudy (amistre64):

F(b)-F(a) = 0 assumes that at a and b, we are in the same place. Either a = b, or we have simply ended were we started.

OpenStudy (anonymous):

Then the location of the bee is still the origin at time T?

OpenStudy (amistre64):

yes, at time T, the bee is simply back in the same place it was at when the time was equal to 0

OpenStudy (anonymous):

But why substitute u in for t for the integration? That's the most confusing part. Does that have any effect on the result?

OpenStudy (amistre64):

0 and T are values that the variable u can take on. the function itself is dependant on the value of u ...

OpenStudy (anonymous):

What is the difference between ∫(0-T) r'(u)du = 0 and ∫(0-T) r'(t)dt = 0 Why did they substitute ?

OpenStudy (amistre64):

hmm, im not absolutely sure, but i think it has something to do with what is called a 'dummy' variable - meant to avoid confusion between the overuse of the letter T

OpenStudy (anonymous):

oh ok .. well thank you !

OpenStudy (amistre64):

good luck